# A Closer Look at Model Collapse: From a Generalization-to-Memorization Perspective

### v1

@@ Abstract
The widespread use of diffusion models has led to an abundance of AI-generated data, raising concerns about model collapse—a phenomenon in which recursive iterations of training on synthetic data lead to performance degradation. Prior work primarily characterizes this collapse via variance shrinkage or distribution shift, but these perspectives miss practical manifestations of model collapse. This paper identifies a transition from generalization to memorization during model collapse in diffusion models, where models increasingly replicate training data instead of generating novel content during iterative training on synthetic samples. This transition is directly driven by the declining entropy of the synthetic training data produced in each training cycle, which serves as a clear indicator of model degradation. Motivated by this insight, we propose an entropy-based data selection strategy to mitigate the transition from generalization to memorization and alleviate model collapse. Empirical results show that our approach significantly enhances visual quality and diversity in recursive generation, effectively preventing collapse.
@@ 2 Background
Diffusion Models. For a given data distribution, diffusion models do not directly learn the probability density function (pdf) of the distribution; instead, they define a forward process and a reverse process, and learn the score function utilized in the reverse process. Specifically, the forward process (Ho et al., 2020) progressively adds Gaussian noise to the image, and the conditional distribution of the noisy image is given by: $p(\bm{x_{t}}|\bm{x_{0}})=\mathcal{N}\left(\bm{x_{t}};\sqrt{\bar{\alpha}_{t}}\bm{x_{0}},(1-\bar{\alpha}_{t})\bm{I}\right)$, where $\bar{\alpha}_{t}$ is the scale schedule, $\bm{x_{0}}$ is the clean image drawn from $P_{0}$, and $\bm{x_{t}}$ is the noisy image. This forward can also be described as a stochastic differential equation (SDE) (Song and Ermon, 2019): $d\bm{x}=f(\bm{x},t)dt+g(t)d\bm{w}$, where $f(\cdot,t):\mathbb{R}^{d}\to\mathbb{R}^{d}$ denotes the vector-valued drift coefficient, $g(t)\in\mathbb{R}$ is the diffusion coefficient, and $w$ is a standard Brownian motion. This SDE has a corresponding reverse SDE as $d\bm{x}=[f(\bm{x},t)-g^{2}(t)\nabla_{\bm{x}}\log p_{t}(\bm{x})]dt+g(t)d\bm{w}$, where $dt$ represents a negative infinitesimal time step, driving the process from $t=T$ to $t=0$. The reverse SDE enables us to gradually convert a Gaussian noise to a clean image $\bm{x}\sim P_{0}$.
@@
and can be efficiently optimized with score match methods such as denoising score matching (Vincent, 2011).
Self-Consuming Loop. Following the standard setting of model collapse (Alemohammad et al., 2024b; Dohmatob et al., 2024b; Kazdan et al., 2024; Dohmatob et al., 2024d, a; Alemohammad et al., 2024a; Feng et al., 2024), we denote the training dataset at the $n$-th iteration as $\mathcal{D}_{n}$. Let $\mathcal{A}(\cdot)$ denote the training algorithm that takes $\mathcal{D}_{n}$ as input and outputs a diffusion model characterized by the distribution $P_{n}$, i.e., $P_{n}=\mathcal{A}(\mathcal{D}_{n})$. In this work, we train the diffusion model from scratch at each iteration. Subsequently, the diffusion model generates a synthetic dataset of size $N_{n}$, denoted by $\mathcal{G}_{n}\sim P_{n}^{N_{n}}$, which is used in subsequent iterations.
@@ 3 Model Collapses from Generalization to Memorization
In this section, we empirically demonstrate the transition from generalization to memorization that occurs over recursive iterations, and investigate the underlying factors driving this transition. All experiments in this section are conducted on the CIFAR-10 dataset (Krizhevsky et al., 2009) using a UNet-based DDPM model (Ho et al., 2020), under the replace paradigm, where each model is trained solely on samples generated by the model from the previous iteration.
Generalization Score. To quantify generalization ability, we adopt the generalization score (Yoon et al., 2023; Alaa et al., 2022; Zhang et al., 2024), defined as the average distance between each generated image and its nearest training image:
@@
Remark: Zhang et al. (2024) measure generalizability as the probability that the similarity between a generated image and its nearest training sample exceeds a threshold. Alaa et al. (2022) assesse memorization via a hypothesis test. While definitions of generalizability vary across studies, they are fundamentally similar, relying on nearest neighbor distances.
Highlight of Observations: The generated data progressively collapses into numerous compact local clusters over model collapse iterations, as evidenced by both the sharp decline in entropy over iterations and visualizations. This localized concentration of data points then facilitates memorization in subsequent models, reducing their ability to generate novel images. Our claim is supported by the following three findings.
Finding I: Generalization-to-Memorization Transition
@@
Finding II: The Entropy of the Training Set Shrinks Sharply over Iterations.
@@
Estimation. However, the density function $f$ of the image distribution is unknown. We use the following Kozachenko-Leonenko (KL) estimator proposed in a well-known paper (Kozachenko, 1987) to empirically estimate $H(\bm{X})$ from a finite set of i.i.d. samples $\mathcal{D}$} drawn from the distribution $P$:
@@
Finding III: The Correlation between Entropy and Generalization Score
We verify that the generalization score of the trained model is strongly correlated with the entropy of the training dataset. The similar collapsing trends of entropy and the generalization score motivate a deeper investigation into their relationship. In Figure˜4(a), we present a scatter plot of entropy versus generalization score across different dataset sizes and successive iterations, with the y-axis shown on a logarithmic scale. Notably, the entropy of the training dataset exhibits a significant linear relationship with the logarithm of the generalization score. The Pearson correlation coefficient is $0.91$ with a $p$-value near zero, quantitatively confirming the strength of this correlation. Furthermore, the scatter points corresponding to different dataset sizes are all approximately aligned along a single line, suggesting the generality of the relationship between entropy and generalization. Specifically, training datasets with higher entropy consistently yield better generalization in the trained model. We also observe that larger datasets typically have higher entropy and result in better generalization, aligning with the conclusion in Zhang et al. (2024) regarding the connection between dataset size and generalization. For comparison, Figure˜4(b) shows the scatter plot of variance versus generalization score. The correlation appears substantially weaker than in Figure˜4(a), suggesting that variance in the training dataset may not directly influence the generalization performance of the trained model.
Conclusion. Findings I–III collectively indicate that the generated data gradually collapses into compact clusters, as shown by declining entropy. This concentration promotes memorization in the model of the next iteration and reduces their ability to generalize.
@@ 4 Mitigating Model Collapse with Data Selection Methods
Method I: Greedy Selection. The procedure iteratively constructs a subset ${\mathcal{D}}\subset\mathcal{S}$ of size $n$ by adding the farthest point one at a time as follows:
@@
Methods II: Threshold Decay Filter. The procedure constructs a subset ${\mathcal{D}}\subset\mathcal{S}$ of size $N$ by iteratively selecting samples that are sufficiently distant from the current set $\mathcal{D}$}. The algorithm proceeds as follows:
@@ 5 Experiments
Datasets. We conduct experiments on three widely used image generation benchmarks. CIFAR-10 (Krizhevsky et al., 2009) consists of $32\times 32$ color images in $10$ classes. Due to computational constraints, we use a subset of 32,76832{,}768 training images. Our goal is not to achieve state-of-the-art FID among large diffusion models but rather to demonstrate that our method mitigates memorization in the self-consuming loop. As shown in Section˜3, this subset is sufficient to observe the transition from generalization to memorization. We also conduct experiments on subsets of FFHQ (Jain et al., 2004), downsampled to $32\times 32$ resolution, and MNIST (LeCun et al., 1998), using 8,1928{,}192 and 12,00012{,}000 training images, respectively.
@@
Model. For CIFAR-10 and FFHQ, we employ a UNet-based backbone (Ronneberger et al., 2015) designed for $32\times 32$ RGB images, which predicts noise residuals. The network contains approximately 16M parameters. For MNIST, we use a similar UNet-based architecture adapted for single-channel inputs, with a total of 19M parameters. Detailed network configurations are provided in the Appendix.
@@
Implementation. For iterative training and sampling, our implementation is based on the Hugging Face Diffusers codebase (von Platen et al., 2022) of DDPM. We use a mixed-precision training of FP16 to train the models. We adopt an Adam optimizer with a learning rate of $10^{-4}$ and a weight decay of $10^{-6}$. The batch size is $128$. A $1000$-step denoising process is used, with all other hyperparameters set to their default values. For Threshold Decay Filter, we use an initial threshold of $60$ and a decay rate of $0.95$. We show in the Appendix that our method is robust in a wide range of hyperparameters.
@@
Evaluation Metrics. We use the generalization score and entropy to evaluate the effectiveness of our method in mitigating memorization. Additionally, we adopt the Fréchet Inception Distance (FID) (Heusel et al., 2017) as a metric to quantify the distributional divergence between generated images and real images.
@@
Results. To evaluate the efficacy of our selection methods in mitigating memorization within the self-consuming loop, we compare the generalization scores of two vanilla paradigms with those augmented by Greedy Selection or Threshold Decay Filter. As shown in Figure˜5, Greedy Selection consistently improves generalization scores across all datasets and paradigms and is particularly effective in the accumulate-subsample setting. Importantly, our selection methods mitigate memorization without compromising FID performance; in fact, they can even slow down FID degradation. Figure˜6 reports the FID of generated images over successive iterations. Under the accumulate-subsample paradigm, both methods yield notable improvements in FID, with Greedy Selection outperforming Threshold Decay Filter. For example, the vanilla accumulate paradigm reaches an FID of $75.7$ at iteration $8$, whereas Greedy Selection significantly reduces it to $44.7$. On the FFHQ dataset under the replace paradigm, however, Threshold Decay Filter performs better, suggesting that adaptive selection strength may be beneficial in certain cases.
@@
Analysis for the Improvement. We present the estimated entropy of the training datasets in Figure˜7. As shown in the figure, our selection methods effectively increase the entropy of the training data at each iteration, consistent with their design. These more diverse, higher-entropy datasets subsequently enable the next-iteration model to generalize better, as evidenced by the improved generalization scores in Figure˜5.
@@
Diversity Improvement on Classifier-Free Diffusion Guidance (CFG) (Ho and Salimans, 2022). We further validate the effectiveness of our methods in the CFG setting and show that they substantially improve the diversity of generated images. CFG is a widely used conditional generation technique that consistently enhances perceptual quality but often sacrifices diversity (Chidambaram et al., 2024; Ho and Salimans, 2022; Li et al., 2025a). Prior work by Yoon et al. (2024) identifies the CFG scale as a key factor influencing the rate of model collapse and suggests that setting a moderate scale can help mitigate model collapse. Experimentally, we observe that the CFG generation can indeed generate clearer images than the unconditional baseline, as compared in Figures˜9(a) and 9(b). However, the diversity of generated images collapses rapidly even with a modest guidance scale, with samples within each class soon becoming nearly identical. In contrast, Figures˜9(c) and 9(d) demonstrate that augmenting CFG with our data selection methods significantly improves image diversity and yields substantially lower FID scores compared to the vanilla CFG paradigm. These results demonstrate that our approach effectively mitigates the diversity loss of CFG while preserving its quality advantage throughout the self-consuming loop.
@@ Appendix D Additional Results
D.1 Generated Images over Iterations
We provide additional generated images of the trained model across iterations and dataset sizes in a grid format.
Figures˜12, 13, 14, 15, 16 and 17 present the generated images of the model trained on CIFAR-10 with various dataset sizes across successive iterations. These images are generated from the vanilla replace paradigm.
As shown in Figures˜12 and 13, the diffusion models tend to memorize small training datasets throughout the recursive process, exactly copying training images during generation. Because the generated samples are nearly identical to the training data, image quality does not noticeably degrade. However, duplicated generations emerge in later iterations, and diversity declines as the model gradually loses coverage over parts of the original images.

### v2

@@ Abstract
The widespread use of diffusion models has led to an abundance of AI-generated data, raising concerns about model collapse—a phenomenon in which recursive iterations of training on synthetic data lead to performance degradation. Prior work primarily characterizes this collapse via variance shrinkage or distribution shift, but these perspectives miss practical manifestations of model collapse. This paper identifies a transition from generalization to memorization during model collapse in diffusion models, where models increasingly replicate training data instead of generating novel content during iterative training on synthetic samples. This transition is directly driven by the declining entropy of the synthetic training data produced in each training cycle, which serves as a clear indicator of model degradation. Motivated by this insight, we propose an entropy-based data selection strategy to mitigate the transition from generalization to memorization and alleviate model collapse. Empirical results show that our approach significantly enhances visual quality and diversity in recursive generation, effectively preventing collapse. The source code is available at https://github.com/shilianghe007/Model_Collapse.git
@@ 2 Background
Diffusion Models.
For a given data distribution, diffusion models do not directly learn the probability density function (pdf) of the distribution; instead, they define a forward process and a reverse process, and learn the score function utilized in the reverse process. Specifically, the forward process (Ho et al., 2020) progressively adds Gaussian noise to the image, and the conditional distribution of the noisy image is given by: $p(\bm{x_{t}}|\bm{x_{0}})=\mathcal{N}\left(\bm{x_{t}};\sqrt{\bar{\alpha}_{t}}\bm{x_{0}},(1-\bar{\alpha}_{t})\bm{I}\right)$, where $\bar{\alpha}_{t}$ is the scale schedule, $\bm{x_{0}}$ is the clean image drawn from $P_{0}$, and $\bm{x_{t}}$ is the noisy image. This forward can also be described as a stochastic differential equation (SDE) (Song and Ermon, 2019): $d\bm{x}=f(\bm{x},t)dt+g(t)d\bm{w}$, where $f(\cdot,t):\mathbb{R}^{d}\to\mathbb{R}^{d}$ denotes the vector-valued drift coefficient, $g(t)\in\mathbb{R}$ is the diffusion coefficient, and $w$ is a standard Brownian motion. This SDE has a corresponding reverse SDE as $d\bm{x}=[f(\bm{x},t)-g^{2}(t)\nabla_{\bm{x}}\log p_{t}(\bm{x})]dt+g(t)d\bm{w}$, where $dt$ represents a negative infinitesimal time step, driving the process from $t=T$ to $t=0$. The reverse SDE enables us to gradually convert a Gaussian noise to a clean image $\bm{x}\sim P_{0}$.
@@
and can be efficiently optimized with score matching methods such as denoising score matching (Vincent, 2011).
Self-Consuming Loop.
Following the standard setting of model collapse (Alemohammad et al., 2024b; Dohmatob et al., 2024b; Kazdan et al., 2024; Dohmatob et al., 2024d, a; Alemohammad et al., 2024a; Feng et al., 2024), we denote the training dataset at the $n$-th iteration as $\mathcal{D}_{n}$. Let $\mathcal{A}(\cdot)$ denote the training algorithm that takes $\mathcal{D}_{n}$ as input and outputs a diffusion model characterized by the distribution $P_{n}$, i.e., $P_{n}=\mathcal{A}(\mathcal{D}_{n})$. In this work, we train the diffusion model from scratch at each iteration. Subsequently, the diffusion model generates a synthetic dataset of size $N_{n}$, denoted by $\mathcal{G}_{n}\sim P_{n}^{N_{n}}$, which is used in subsequent iterations.
@@ 3 Model Collapses from Generalization to Memorization
In this section, we empirically demonstrate the transition from generalization to memorization that occurs over recursive iterations, and investigate the underlying factors driving this transition. All experiments in this section are conducted on the CIFAR-10 dataset (Krizhevsky et al., 2009) using a UNet-based DDPM model (Ho et al., 2020), under the replace paradigm, where each model is trained solely on samples generated by the model from the previous iteration. However, in Appendix˜D, we extend the explorations to other datasets and paradigms, where the conclusion remains valid.
Generalization Score.
To quantify generalization ability, we adopt the generalization score (Yoon et al., 2023; Alaa et al., 2022; Zhang et al., 2024), defined as the average distance between each generated image and its nearest training image:
@@
Remark:
Zhang et al. (2024) measure generalizability as the probability that the similarity between a generated image and its nearest training sample exceeds a threshold. Alaa et al. (2022) assesse memorization via a hypothesis test. While definitions of generalizability vary across studies, they are fundamentally similar, relying on nearest neighbor distances.
Highlight of Observations:
The generated data progressively collapses into numerous compact local clusters over model collapse iterations, as evidenced by both the sharp decline in entropy over iterations and visualizations. This localized concentration of data points then facilitates memorization in subsequent models, reducing their ability to generate novel images. Our claim is supported by the following three findings.
3.1 Finding I: Generalization-to-Memorization Transition
@@
3.2 Finding II: The Entropy of the Training Set Shrinks Sharply over Iterations
@@
Estimation.
However, the density function $f$ of the image distribution is unknown. We use the following Kozachenko-Leonenko (KL) estimator proposed in a well-known paper (Kozachenko, 1987) to empirically estimate $H(\bm{X})$ from a finite set of i.i.d. samples $\mathcal{D}$} drawn from the distribution $P$:
@@
3.3 Finding III: The Correlation between Entropy and Generalization Score
We verify that the generalization score of the trained model is strongly correlated with the entropy of the training dataset. The similar collapsing trends of entropy and the generalization score motivate a deeper investigation into their relationship. In Figure˜4(a), we present a scatter plot of entropy versus generalization score across different dataset sizes and successive iterations, with the y-axis shown on a logarithmic scale. Notably, the entropy of the training dataset exhibits a significant linear relationship with the logarithm of the generalization score. The Pearson correlation coefficient is $0.91$ with a $p$-value near zero, quantitatively confirming the strength of this correlation. Furthermore, the scatter points corresponding to different dataset sizes are all approximately aligned along a single line, suggesting a dataset size-independent relationship between entropy and generalization. Specifically, training datasets with higher entropy consistently yield better generalization in the trained model. We also observe that larger datasets typically have higher entropy and result in better generalization, aligning with the conclusion in Zhang et al. (2024) regarding the connection between dataset size and generalization. For comparison, Figure˜4(b) shows the scatter plot of variance versus generalization score. The correlation appears substantially weaker than in Figure˜4(a), suggesting that variance in the training dataset may not directly influence the generalization performance of the trained model.
Conclusion.
Findings I–III collectively indicate that the generated data gradually collapses into compact clusters, as shown by declining entropy. This concentration promotes memorization in the model of the next iteration and reduces its ability to generalize. In Appendix˜D, we further extend these explorations to the FFHQ dataset and accumulate paradigm, demonstrating a more robust relation.
@@ 4 Mitigating Model Collapse with Data Selection Methods
Method I: Greedy Selection.
The procedure iteratively constructs a subset ${\mathcal{D}}\subset\mathcal{S}$ of size $n$ by adding the farthest point one at a time as follows:
@@
Methods II: Threshold Decay Filter.
The procedure constructs a subset ${\mathcal{D}}\subset\mathcal{S}$ of size $N$ by iteratively selecting samples that are sufficiently distant from the current set $\mathcal{D}$}. The algorithm proceeds as follows:
@@ 5 Experiments
We conduct experiments on three widely used image generation benchmarks. CIFAR-10 (Krizhevsky et al., 2009) consists of $32\times 32$ color images in $10$ classes. Due to computational constraints, we use a subset of 32,76832{,}768 training images. Our goal is not to achieve state-of-the-art FID among large diffusion models but rather to demonstrate that our method mitigates memorization in the self-consuming loop. As shown in Section˜3, this subset is sufficient to observe the transition from generalization to memorization. We also conduct experiments on subsets of FFHQ (Jain et al., 2004), downsampled to $32\times 32$ resolution, and MNIST (LeCun et al., 1998), using 8,1928{,}192 and 12,00012{,}000 training images, respectively.
Model.
@@
For CIFAR-10 and FFHQ, we employ a UNet-based backbone (Ronneberger et al., 2015) designed for $32\times 32$ RGB images, which predicts noise residuals. The network contains approximately 16M parameters. For MNIST, we use a similar UNet-based architecture adapted for single-channel inputs, with a total of 19M parameters. Detailed network configurations are provided in the Appendix.
Implementation.
@@
For iterative training and sampling, our implementation is based on the Hugging Face Diffusers codebase (von Platen et al., 2022) of DDPM. We use a mixed-precision training of FP16 to train the models. We adopt an Adam optimizer with a learning rate of $10^{-4}$ and a weight decay of $10^{-6}$. The batch size is $128$. A $1000$-step denoising process is used, with all other hyperparameters set to their default values. For Threshold Decay Filter, we use an initial threshold of $60$ and a decay rate of $0.95$. We show in the Appendix that our method is robust in a wide range of hyperparameters.
Evaluation Metrics.
@@
We use the generalization score and entropy to evaluate the effectiveness of our method in mitigating memorization. Additionally, we adopt the Fréchet Inception Distance (FID) (Heusel et al., 2017) as a metric to quantify the distributional divergence between generated images and real images.
Results.
@@
To evaluate the efficacy of our selection methods in mitigating memorization within the self-consuming loop, we compare the generalization scores of two vanilla paradigms with those augmented by Greedy Selection or Threshold Decay Filter. As shown in Figure˜5, Greedy Selection consistently improves generalization scores across all datasets and paradigms and is particularly effective in the accumulate-subsample setting. Importantly, our selection methods mitigate memorization without compromising FID performance; in fact, they can even slow down FID degradation. Figure˜6 reports the FID of generated images over successive iterations. Under the accumulate-subsample paradigm, both methods yield notable improvements in FID, with Greedy Selection outperforming Threshold Decay Filter. For example, the vanilla accumulate paradigm reaches an FID of $75.7$ at iteration $8$, whereas Greedy Selection significantly reduces it to $44.7$. On the FFHQ dataset under the replace paradigm, however, Threshold Decay Filter performs better, suggesting that adaptive selection strength may be beneficial in certain cases.
@@
Analysis for the Improvement.
We present the estimated entropy of the training datasets in Figure˜7. As shown in the figure, our selection methods effectively increase the entropy of the training data at each iteration, consistent with their design. These more diverse, higher-entropy datasets subsequently enable the next-iteration model to generalize better, as evidenced by the improved generalization scores in Figure˜5.
@@
Diversity Improvement on Classifier-Free Diffusion Guidance (CFG) (Ho and Salimans, 2022).
We further validate the effectiveness of our methods in the CFG setting and show that they substantially improve the diversity of generated images. CFG is a widely used conditional generation technique that consistently enhances perceptual quality but often sacrifices diversity (Chidambaram et al., 2024; Ho and Salimans, 2022; Li et al., 2025a). Prior work by Yoon et al. (2024) identifies the CFG scale as a key factor influencing the rate of model collapse and suggests that setting a moderate scale can help mitigate model collapse. Experimentally, we observe that the CFG generation can indeed generate clearer images than the unconditional baseline, as compared in Figures˜9(a) and 9(b). However, the diversity of generated images collapses rapidly even with a modest guidance scale, with samples within each class soon becoming nearly identical. In contrast, Figures˜9(c) and 9(d) demonstrate that augmenting CFG with our data selection methods significantly improves image diversity and yields substantially lower FID scores compared to the vanilla CFG paradigm. These results demonstrate that our approach effectively mitigates the diversity loss of CFG while preserving its quality advantage throughout the self-consuming loop.
@@ B.3 Pseudo-codes for the Algorithms
We present the pseudo-code of the Greedy Selection and Threshold Decay Filter methods in Algorithms˜1 and 2.
@@ C.2 Different Decay Rates
Figure 12: The generalization-to-memorization transition on the accumulate paradigm. Similar to Figure˜2, we visualize the generated images and their nearest neighbors in the training dataset. As the iteration proceeds, the model loses generalization ability, collapsing into the memorization regime.
Figure 13: Scatter plots of the generalization score and estimated entropy of the training data. Each point denotes one iteration of training in the self-consuming loop. In addition to the results in Figure˜4(a), we include more results from FFHQ and on the accumulate paradigm. We use different colors to denote the loops, different shapes (circles and triangles) to represent CIFAR and FFHQ, and solid versus hollow markers to distinguish the replace and accumulate paradigms.
@@ Appendix D Additional Results
D.1 The Generalization-to-Memorization Collapse on Accumulate Paradigm
This subsection presents additional results on the accumulate paradigm and FFHQ dataset as a supplement to the explorations in Section˜3. Specifically, we show in Figure˜12 that the generalization-to-memorization collapse also occurs on the accumulation paradigm. We note that model collapse has different definitions in previous studies. Here, we clarify that our results show the generalization score consistently decreases over early iterations. However, we cannot determine whether the score eventually collapses to zero or converges to a lower bound in the accumulate paradigm, as we only have fewer than 10 iterations—insufficient to draw conclusions about convergence.
D.2 The Robust Relationship between Generalization Score and Entropy
In Figure˜2, we plot the results on CIFAR-10 datasets, demonstrating the dataset size-independent relationship between generalization score and estimated entropy. This subsection further incorporates the results from FFHQ and the accumulate paradigm.
Specifically, we conduct the self-consuming loop for FFHQ with dataset sizes of $8,192$, $16,384$, and $32,768$. Besides, we also include the results from Section˜D.1. Combining those, we show a more robust relationship in Figure˜13, where all the points align around the red dashed line. This result suggests that the entropy of the training dataset is a dataset-independent indicator for the generalization score.
D.3 Generated Images over Iterations
This subsection provides additional generated images of the trained model across iterations and dataset sizes in a grid format. These images are generated from the vanilla replace paradigm.
Figures˜14, 15, 16, 17, 18 and 19 present the generated images of the model trained on CIFAR-10 with various dataset sizes across successive iterations. As shown in Figures˜14 and 15, the diffusion models tend to memorize small training datasets throughout the recursive process, exactly copying training images during generation. Because the generated samples are nearly identical to the training data, image quality does not noticeably degrade. However, duplicated generations emerge in later iterations, and diversity declines as the model gradually loses coverage over parts of the original images.